OpenStudy (anonymous):

13.   A worker takes a random sample of 200 bolts and finds that 30 of them are either too long or too short, thus making them unusable. Estimate the number of unusable bolts in a production of 17,000 bolts.

OpenStudy (anonymous):

its 85?

OpenStudy (anonymous):

yup. now multiply that by 30

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yup :) no prob

jimthompson5910 (jim_thompson5910):

(# of unusable bolts)/(# of bolts total) = (# of unusable bolts)/(# of bolts total) 30/200 = x/17000 30*17000 = 200x 510000 = 200x 510000/200 = x 2550 = x x = 2550 So there are roughly 2550 unusable bolts in a production of 17,000 bolts

OpenStudy (anonymous):

thanks jim to

jimthompson5910 (jim_thompson5910):

yw